Car A is traveling eastward at 85 km/h. Car B is traveling westward with a speed of 65 km/h.

According to classical relativity, the velocity of car B as measured by the driver of car A would be

According to classical relativity, the velocity of car B as measured by the driver of car A would be

20 km/h eastward.
150 km/h westward.
20 km/h westward.
150 km/h eastward.
Answer #1

Vb/a = 85 + 65 = 150 km westward

150 km/h westward.
Answer #2

According to classical relativity, the velocity of car B as measured by the driver of car A would be
Ans: 65+85 = 150 kh/hr westards

According to classical relativity, the velocity of car B as measured by the driver of car A would be
Ans: 65+85 = 150 kh/hr eastward
Answer #3

relative velocity = 65+85 = 150 westwards

when a is stopped car b move with relative velocity in west direction